a disaster relief team consists of engineers and doctors in the ratio of 2:5
  1. if there are 18 engineers, find the number of doctors.
  2. if there are 65 doctors, find the number of engineers.

Grade:7

2 Answers

Ajay Panthagani
7 years ago
let ratio of engineers and doctors be 2x:5x
therefore if there are 18 engineers
then 2x=18
          x=18/2
          x=9
then number of doctors= 5x= 5(9)=45
if there are 65 doctors
then 5x=65
          x=65/5
          x=13
then number of engineers=2x=2(13)=26
Avi
7 years ago
Since the ratio of engineers to doctors is 2:5 we can let the number of engineers to be 2x and number of doctors to be 5x
It is given that there are 18 engineers so according to the above assumption, 2x=18 and x=9
No. of doctors = 5x= 5*9 =45
The second part can be solved similarly as 5x=65 and x=13
So number of engineers = 2x= 2*13=26